Grade 12Physical Chemistry

The volume of 0.1m hcl required to neutralize 20ml of a solution containing 0.106g of na2co3 is

Chaitanya
8 Years agoGrade 12

1 Answer

Abhiraj Pathak
8 Years ago
molar mass of Na2CO3 = 106
moles of Na2CO3 in 0.106g = 0.106/106 = 0.001
Molarity of Na2CO3 in 20mL of solution = 0.001/0.02 = 0.05M
Na2CO3 + 2 HCl = 2NaCl + H2O + CO2
M1V1 = M2V2
V1 = 0.05*20/0.1 = 10mL
Thus, we would require 2 moles of HCl per mole of Na2CO3 or twice the volume of HCl for complete neutralisation
Therefore volume of 0.1m HCl required is 20mL